OpenStudy (anonymous):

if a >0 and b <0 which of the following is true b-a > 0 a+b <0 a+b >0 a-b > 0 ab>0

OpenStudy (anonymous):

I think No.4 is the only correct answer. Let's imagine a is 3 and b is -2. Then, 1. -2-3 > 0 is incorrect. 2. 3+(-2)=1 < 0 also incorrect 3. 3+(-2)=1 > 0 this one is correct in this case. But let's think about it. What if b was -4 now? In that case, a+b would give us -1 which is less than 0, making the premise No.2 true under those numbers. 4. 3-(-2)=5 > 0. It's easy to see this is correct AND will ALWAYS be correct as long as a is a positive number and b is negative. 5. 3 * (-2) > 0 is incorrect, and always will be so if there's a negative and a positive number (that means the result will always be negative). Hope I helped :)

OpenStudy (anonymous):

i dont know